OpenStudy (anonymous):

One number is 4 more than another, and their sum is 60. What is the smaller number? If x = the larger number and y = the smaller number, then which of the following systems could be used to solve the problem? x + y = 60 and -x + y = 4 x + y = 60 and x - y = 4 x + y = 60 and x + y = 4

OpenStudy (anonymous):

@TrojanPoem

OpenStudy (anonymous):

Since x is the larger number so it x=y+4

OpenStudy (anonymous):

So b

OpenStudy (trojanpoem):

x = y + 4 x -y = 4 y + x = 60

OpenStudy (trojanpoem):

as @joyraheb said.

OpenStudy (anonymous):

thank you:)